OpenStudy (anonymous):

solve algebraically for x: x+2/6 = 3/x-1. (/ symbol makes it a fraction)

OpenStudy (anonymous):

x=-5 or x=4

OpenStudy (angela210793):

x+2/6 = 3/x-1 x+1/2=3/x-1 x-3/x=-1-1/2 (x^2-3)/x=-3/2 2*(x^2-3)=-3*x 2x^2-6+3x=0 D=b^2-4ac=36-4*2*3=12>0 X1=(-b-sqrD)/2a X2=(-b+sqrD)2a I guess....

OpenStudy (anonymous):

A little error at the beginning: 2/6 simplifies to 1/3, not 1/2.

OpenStudy (angela210793):

yea..i was wondering why D's sqr wan't a integer...thanks :):):)

OpenStudy (angela210793):

So figgles4 just put 1/3 instead of 1/2 and do what i have done :)

OpenStudy (anonymous):

alright, thank you everyone, i really appreciate all your help:D